# DiverseDream: Diverse Text-to-3D Synthesis with Augmented Text Embedding

### v1

@@ Abstract
Text-to-3D synthesis has recently emerged as a new approach to sampling 3D models by adopting pretrained text-to-image models as guiding visual priors. An intriguing but underexplored problem with existing text-to-3D methods is that 3D models obtained from the sampling-by-optimization procedure tend to have mode collapses, and hence poor diversity in their results. In this paper, we provide an analysis and identify potential causes of such a limited diversity, and then devise a new method that considers the joint generation of different 3D models from the same text prompt, where we propose to use augmented text prompts via textual inversion of reference images to diversify the joint generation. We show that our method leads to improved diversity in text-to-3D synthesis qualitatively and quantitatively.
@@ 1 Introduction
While text-to-3D synthesis has shown interesting results, an intriguing issue is that the generated 3D models remain limited in several aspects including fidelity, diversity, optimization convergence, and scalability. Some efforts have been made to address these problems, ranging from an improved loss function in ProlificDreamer [47], amortized generation over different text prompts in ATT3D [27], to personalized generation in DreamBooth3D [35]. Among these leading efforts to improve text-to-3D synthesis, diversity has been an underexplored aspect, and there is little insight into how diversity is achieved in existing text-to-3D methods.
In this paper, we investigate the diversity of 3D model generation from state-of-the-art text-to-3D methods. We argue that the diversity is affected by the objective function such as SDS [33] and VSD [47] given a text prompt to condition the 3D model generation. Inspired by this observation, we propose to diversify text-to-3D generation results by applying an augmentation to the original text prompt in the text embedding via textual inversion techniques [12, 15]. Our idea is to first sample reference images from a pretrained text-to-image diffusion model and then find the corresponding text features of these images. By using these augmented text features to condition the optimization process to sample 3D models instead of the original text prompt, we show that the diversity of the 3D models can be significantly improved. Our experiments show that our method outperforms the state-of-the-art text-to-3D methods in terms of diversity quantitatively and qualitatively.
@@
  • Extensive experiments and ablation studies that demonstrate the validity and robustness of our method, which is applicable to different text-to-3D methods;
@@ 2 Related Work
Text-to-image synthesis has witnessed tremendous growth in the past years. The first group of methods [41, 19, 3] are built upon Generative Adversarial Networks (GANs), which leverage an actor-critic loss for training their image generators. Despite the capability of GANs to generate fast and realistic images, they are susceptible to mode collapse. An alternative group of approaches is auto-regressive models, for example, DALL-E [36], Parti [52], and MUSE [6] which model the generation of an image patch based on another one, resulting in a sequential generation modeling. A recent notable family of text-to-image synthesis is based on diffusion models, such as Stable Diffusion [38], DALL-E 3 [42], and Imagen[40], which have demonstrated great potential to synthesize high-resolution high-quality images. In this paper, we resort to diffusion-based models, e.g., Stable Diffusion, as the pretrained 2D prior for supervising our 3D model generation.
@@
3D representation is the core of generative 3D models for various 3D tasks, such as novel view synthesis and 3D content creation. Neural Radiance Fields (NeRFs) [29], distinguished for its utilization of volumetric rendering, has gained considerable attraction in learning a 3D scene representation under solely 2D image supervision. Despite NeRF’s pervasive application [46] in both 3D reconstruction and generation domains, its optimization process has been identified as time-intensive [13]. There is a vast amount of work addressing the speed convergence issue and improving view synthesis performance using hybrid neural scene representations such as voxel grids [44, 51], hash-grid [30], tri-planes [4, 7] and Gaussian splatting [21]. Among those hybrid scene representations, hash-grids [30] has been extensively used for the text-to-3D task [33, 47] due to its fast training, superior performance compared to vanilla NeRFs, and absence of bias phenomena observed in the tri-planes. This paper also utilizes hash grids to learn a set of diverse 3D scene representations from a single text prompt via our proposed textual score distillation loss.
Image-to-3D generation is an important direction of conditional generative 3D models. Early work on single view synthesis [49, 31] leverages a differentiable neural renderer to warp the input view and then refine it to perform novel view synthesis. However, such approaches are limited by the distance between the input and target pose and it is not trivial to produce a full 360°reconstruction given a single input. There has been significant progress with the emerging generative models. Zero-1-to-3 [25] pioneers open-world single-image-to-3D conversion through zero-shot novel view synthesis. However, the geometric inconsistency in its generated images has yet to bridge the gap between multi-view images and 3D scenes. Recent works such as One-2-3-45 [24], SyncDreamer [26], and Consistent123 [48] build extra geometry-constraint layers upon Zero-1-to-3 to obtain more 3D-consistent results. However, training image-to-3D models usually requires 3D model datasets such as ShapeNet [5], CO3D [37], or large-scale multiview datasets such as Objaverse [10] or Objaverse-XL [9]. Our approach, on the other hand, is only based on a pretrained text-to-image model for supervision, which is more accessible.
Text-to-3D generation. Recent advances in this topic have demonstrated outstanding potential for generating complex 3D models based solely on 2D priors by utilizing pre-trained text-to-image models such as Stable Diffusion [38]. DreamField [18] is a notable early work in this field that employs CLIP [34] to minimize the disparities between the rendered image and the input text across a predetermined set of random perspectives. However, the quality of 3D models generated by DreamField is often compromised due to the limited ability of the CLIP model to capture the rendered image’s semantic features fully. In response to this limitation, DreamFusion [33] proposes a substitution of the CLIP loss with the Score Distillation Sampling (SDS) loss and introduces an efficient method for calculating the gradients to learn a neural radiance field for 3D synthesis. Unfortunately, DreamFusion tends to generate objects with oversmooth surfaces and saturated colors. Several subsequent methods aim to improve upon DreamFusion in many directions: higher resolution [23], richer appearance [8], faster speed [45], photorealistic appearance [20, 47], single view [28, 11, 50]. Despite such rapid development, limited efforts have been spent to investigate the diversity aspect of text-to-3D synthesis, which motivates our work.
@@
Textual inversion. Recent text-to-image diffusion models [38, 42] manage to render high-quality 2D images, but they can not guarantee to preserve the shape and identity of the subject or “personalization”. There exist inversion methods to maintain the identity of the subject in the reference images such as Textual Inversion [12] and DreamBooth [39]. They introduce a virtual token whose corresponding embedding can be optimized and manipulate the generated 2D images. Based on Textual Inversion, HiPer inversion [15] further enhances the inversion capability by using only one reference image and storing the object identity in some optimized textual tokens of the text prompt. This key observation motivates us to apply textual inversion to diversify the generated 3D content.
@@ 3 Background
$\nabla_{\theta}\mathcal{L}_{\text{SDS}}=\mathbb{E}_{t,\epsilon,c}\left[\omega(t)(\epsilon_{\text{SD}}(x_{t},t,y)-\epsilon)\frac{\partial g(\theta,c)}{\partial\theta}\right],$ (1)
where $\omega(t)$ is a time-dependent weighting function, $\epsilon_{\text{SD}}$ is the predicted noise of SD given the noisy input image $x_{t}=\alpha_{t}x+\sigma_{t}\epsilon$ created by adding Gaussian noise $\epsilon$ to the rendered image $x$ at timestep $t$ with noise scheduling coefficients $\alpha_{t},\sigma_{t}$. However, the SDS loss often suffers from over-saturation, over-smoothing, and low-diversity issues as empirically analyzed in [47].
The particular issue of low diversity in SDS can be empirically observed in that the optimization from different runs tends to give similar results.
Therefore, we advocate the use of the more sophisticated variational score distillation loss [47] for our exploration of the diversity of text-to-3D synthesis, which we briefly describe below.
Variational score distillation (VSD): ProlificDreamer [47] addresses the aforementioned limitations of DreamFusion by proposing a variational form of score distillation. Their VSD loss attempts to address the low-diversity issue by modeling the distribution of 3D model output from a single text prompt as $\mu(\theta|y)$. Note that SDS is a special case of VSD, when $\mu(\theta|y)$ reduces to a Dirac distribution $\delta(\theta-\theta^{1})$, having only a single 3D representation $\theta^{1}$ for each text prompt.
@@
$\nabla_{\theta_{i}}\mathcal{L}_{\text{VSD}}\triangleq\\ \mathbb{E}_{t,\epsilon,c}\left[\omega(t)(\epsilon_{\text{SD}}(x_{t}^{i},t,y)-\epsilon_{\phi}(x_{t}^{i},t,c,y))\frac{\partial g(\theta_{i},c)}{\partial\theta_{i}}\right],$ (2)
@@
$\min_{\phi}\mathbb{E}_{t,\epsilon,c}\left[\|\omega(t)(\epsilon_{\phi}(x_{t},t,c,y)-\epsilon)\|^{2}_{2}\right].$ (3)
@@
3:     Optimize $h_{i}$ given $x^{r}_{i}$ following Eq. (4) to obtain $h_{i}^{*}$.
@@
5:     Update $\theta_{i}$ following Eq. (6).
6:     Update $\phi$ following Eq. (7).
@@
4 Our Approach
Inspired by textual inversion methods [12, 15] for 2D object personalization, we aim to condition text-to-3D synthesis such that the per-particle difference in $\epsilon_{\phi}(x_{t}^{i},t,c,y^{\prime}_{i})$ is boosted by using different and distinct prompt $y^{\prime}_{i}$ for each particle $\theta_{i}$.
Particularly, we consider the distribution of 3D representations from a given text prompt $y$. Our goal is to diversify the sampling for this distribution so that the sampled representations align with the given text prompt while having significant differences in terms of geometry and/or appearance. Our key idea is to diversify the 3D representations by diversifying their language descriptions. We observe that an object can be described in multiple valid ways, implying the existence of various text prompts that correspond to the input prompt $y$. By sampling these valid text prompts, denoted as $y_{i}^{\prime}\sim p(y^{\prime}|y)$, we can distill the 3D models using these prompts, thereby introducing diversity into the model. To this end, we devise a new approach that leverages HiPer textual inversion [15] to enhance the resulting diversity.
Overview. Our method has two stages: HiPer tokens inversion and textual score distillation. In the first stage, we sample a reference image $x^{r}_{i}$ for each particle and find an optimized HiPer token $h^{*}_{i}$ so that the reference image can be faithfully reconstructed by the pretrained text-to-image diffusion model with the prompt $[y;h^{*}_{i}]$, where [;][;][ ; ] is the concatenation operation. In the second stage, multiple particles $\theta_{i}$ are jointly optimized along with a new shared domain adapter $\phi$ which is also encoded as a learnable token to form the augmented text prompt $[y;h^{*}_{i};\phi]$ for each particle. The algorithm is shown in Alg. 1.
@@ 4.1 HiPer tokens inversion
Figure 3: Optimization progression of VSD (upper) vs ours (lower). TSD with less # learnable parameters $\phi$ converges faster than VSD.
@@ 4.2 Textual score distillation (TSD)
$\nabla_{\theta_{i}}\mathcal{L}_{\text{TSD}}\triangleq$ (6)
$\mathbb{E}_{t,\epsilon,c}\left[\omega(t)(\epsilon_{\text{SD}}(x_{t}^{i},t,y_{i})-\epsilon_{\text{SD}}(x_{t}^{i},t,y^{\prime}_{i}))\frac{\partial g(\theta_{i},c)}{\partial\theta_{i}}\right].$
@@
In Fig. 2 (Right), we show how we train the sampled NeRF model $\theta_{i}$ and shared token $\phi$ using the proposed TSD (equation 6) and MSE (equation 7) losses respectively. As can be seen in the Fig. 3, our method can produce higher quality samples than those produced by VSD [47] given the same amount of optimization time.
@@ 5 Experiments
Metrics. The common metrics for generative models such as FID [16] do not separately measure fidelity and diversity. Inspired by [1], we proposed to use a modified version of Inception Quality (IQ) and Inception Variance (IV) to measure the quality and diversity of our models. Our IQ and IV are formulated as follows:
@@
We also propose the Cosine Sim metric to quantify the diversity of our particles. Specifically, for a given set of $K$ particles, we render the same view for each particle. The rendered images are then fed through a feature extractor (e.g., as DINO [53]) to obtain feature vectors. We then calculate the cosine similarity for these feature vectors across (${K\choose 2}$ pairs and take the average. To ensure robustness, we also average the results over 120 rendered views.
IQ normal-$\downarrow$ IV normal-$\uparrow$ Cosine Sim normal-$\downarrow$
SDS [33] 3.4692 4.2915 0.7465
VSD [47] 3.1716 4.2734 0.5383
TSD (ours) 3.5132 4.7181 0.5024
Table 1: Quantitative comparison with SOTA methods.
Implementation details. Our method is implemented with threestudio, an open-source framework for text-to-3D synthesis [14]. We use $K=6$ particles for both VSD and our framework, TSD. The training of all experiments is conducted for 50K iterations. We use $L_{2}=5$ as HiPer tokens optimized in 1.4K iterations. Also, we use $L_{3}=8$ as shared learnable tokens $\phi$ and trained in 50K iterations. For the camera embedding, we use the same implementation as threestudio [14]. Regarding the resolution, we train each particle at $256\times 256$ for all methods. We will release our code upon publication.
Baselines. We compare our method with two prominent text-to-3D methods including DreamFusion [33] and ProlificDreamer [47]. Note that we do not compare to other variants such as Magic3D [23] or Fantasia3D [8] since these methods address different issues of SDS, which is orthogonal to our method which focuses on diversity. For validation, we select a set of 10 text prompts from DreamFusion [33] and 10 text prompts generated randomly from ChatGPT [32].
@@
5.1 Comparison with SOTA

### v2

# Diverse Text-to-3D Synthesis with Augmented Text Embedding

@@ Abstract
Text-to-3D synthesis has recently emerged as a new approach to sampling 3D models by adopting pretrained text-to-image models as guiding visual priors. An intriguing but underexplored problem with existing text-to-3D methods is that 3D models obtained from the sampling-by-optimization procedure tend to have mode collapses, and hence poor diversity in their results. In this paper, we provide an analysis and identify potential causes of such a limited diversity, which motivates us to devise a new method that considers the joint generation of different 3D models from the same text prompt. We propose to use augmented text prompts via textual inversion of reference images to diversify the joint generation. We show that our method leads to improved diversity in text-to-3D synthesis qualitatively and quantitatively. Project page: https://diversedream.github.io/
Keywords:
Text-to-3D 3D computer vision Generative models
**footnotetext: These authors contributed equally to this work
@@ 1 Introduction
While text-to-3D synthesis has shown promises, challenges persist in fidelity, diversity, convergence, and scalability of generated models. Efforts to address these issues include enhanced loss functions like ProlificDreamer [57], generalized across prompts in ATT3D [33], ATOM [39], ET3D [9], and personalized generation in DreamBooth3D [42]. Diversity, however, remains largely unexplored in current text-to-3D methods, with limited insight into its mechanisms.
In this paper, we explore methods to enhance the diversity of 3D model generation in text-to-3D systems. We posit that the diversity of model outputs is influenced by the objective function used, such as SDS [38] and VSD [57], when conditioning 3D model generation on a text prompt. Motivated by this insight, we propose a method to diversify text-to-3D generation results by augmenting the original text prompt through textual inversion techniques [11, 14]. Our approach involves sampling reference images from a pretrained text-to-image diffusion model and extracting the corresponding text features via Textual Inversion. These text features are then combined with the features of the original text prompts to guide the optimization process for sampling 3D models. Experimental results (Fig. 1) demonstrate a significant improvement in the diversity of generated 3D models compared to state-of-the-art methods, both quantitatively and qualitatively.
@@
  • Extensive experiments and ablation studies to demonstrate the validity and robustness of our method, which is applicable to different text-to-3D methods.
@@ 2 Related Work
Text-to-image synthesis has seen significant advancements, with methods relying on Generative Adversarial Networks (GANs) [47, 23, 3] and auto-regressive models like DALL-E [43], Parti [61], and MUSE [6]. While GANs offer fast and realistic image generation, they are prone to mode collapse. Recently, diffusion models such as Stable Diffusion [44], DALL-E 3 [50], and Imagen[46] have shown promise in synthesizing high-quality images. In this study, we utilize diffusion-based models, particularly Stable Diffusion, as the pretrained 2D prior to supervise our 3D model generation.
@@
3D representation serves as the foundation for various 3D tasks like novel view synthesis and content creation. Neural Radiance Fields (NeRFs) [34] have gained traction for their volumetric rendering approach, learning 3D scenes from 2D images alone. Despite NeRF’s widespread use [55], its optimization process is time-intensive [12]. To address this, researchers have explored hybrid scene representations like voxel grids [52, 60], hash-grids [35], tri-planes [4, 7], and Gaussian splatting [25], aiming to improve speed and view synthesis performance. Among these, hash-grids [35] are favored for text-to-3D tasks [38, 57] due to their fast training and superior performance compared to NeRFs. In this paper, we leverage hash grids to learn diverse 3D scene representations from a single text prompt using our proposed textual score distillation loss.
Image-to-3D generation is a crucial aspect of conditional generative 3D models. Early methods like SynSin [59] and Free View Synthesis [36] rely on differentiable neural renderers for single view synthesis but are limited by pose distances and struggle with full 360°reconstructions from a single input. Recent advancements have seen models like Zero-1-to-3 [31] pioneering open-world single-image-to-3D conversion through zero-shot novel view synthesis, yet face challenges with geometric consistency. Works such as One-2-3-45 [30], SyncDreamer [32], LRM [17], LGM [53], and Consistent123 [58] address this by adding geometry-constraint layers to improve consistency. However, these methods typically require extensive 3D model datasets like ShapeNet [5] or large-scale multiview datasets like Objaverse [10] for training. In contrast, our approach solely relies on a pretrained text-to-image model for supervision, making it more accessible.
Text-to-3D generation has seen remarkable progress recently, leveraging pre-trained text-to-image models like Stable Diffusion [44]. Early works like DreamField [21] use CLIP [41] to align rendered images with input text but often compromise on model quality due to CLIP’s limited semantic feature capture. DreamFusion [38] substitutes CLIP loss with Score Distillation Sampling (SDS) and introduces efficient gradient calculation for neural radiance field learning. However, it tends to produce oversmooth surfaces and saturated colors. Subsequent methods aim to address these limitations by enhancing resolution [28], appearance [8, 64, 24, 62], geometry [49, 27, 19, 40, 48], speed [54, 20, 27], and photorealism [24, 57, 26, 64]. Despite this progress, diversity in text-to-3D synthesis remains underexplored, motivating our work.
@@
Textual inversion. While recent text-to-image diffusion models like Stable Diffusion [44] and DALL-E 3 [50] produce high-quality 2D images, they may not preserve the subject’s shape or identity, known as “personalization”. Techniques such as Textual Inversion [11] and DreamBooth [45] aim to maintain subject identity in reference images by introducing a virtual token whose embedding can be optimized to manipulate the generated images. HiPer inversion [14], building upon Textual Inversion, enhances inversion by using a single reference image and optimizing textual tokens in the text prompt to store object identity. Inspired by this, we apply textual inversion to diversify generated 3D content.
@@ 3 Background
$\nabla_{\theta}\mathcal{L}_{\text{SDS}}\triangleq\mathbb{E}_{t,\epsilon,c}\left[\omega(t)(\epsilon_{\text{SD}}(x_{t},t,y)-\epsilon)\frac{\partial g(\theta,c)}{\partial\theta}\right],$ (1)
where $\omega(t)$ is a time-dependent weighting function, $\epsilon_{\text{SD}}$ is the predicted noise of SD given the noisy input image $x_{t}=\alpha_{t}x+\sigma_{t}\epsilon$ created by adding Gaussian noise $\epsilon$ to the rendered image $x$ at timestep $t$ with noise scheduling coefficients $\alpha_{t},\sigma_{t}$. However, the SDS loss often suffers from over-saturation, over-smoothing, and low-diversity issues as empirically analyzed in [57]. The low diversity issue in SDS becomes apparent when multiple runs yield similar results empirically. Therefore, we advocate the use of the more sophisticated variational score distillation loss [57] for our exploration of the diversity of text-to-3D synthesis, which we briefly describe below.
Variational score distillation (VSD): ProlificDreamer [57] mitigates the limitations of DreamFusion by introducing a variational form of score distillation. Their VSD loss aims to tackle the low-diversity issue by modeling the distribution $\mu$ of 3D models $\theta$ generated from a single text prompt $y$ as $\mu(\theta|y)$. It’s worth noting that SDS is a special case of VSD, where $\mu(\theta|y)$ simplifies to a Dirac distribution $\delta(\theta-\theta^{1})$, resulting in only a single 3D model $\theta^{1}$ for each text prompt.
@@
$\nabla_{\theta_{i}}\mathcal{L}_{\text{VSD}}\triangleq\mathbb{E}_{t,\epsilon,c}\left[\omega(t)(\epsilon_{\text{SD}}(x_{t}^{i},t,y)-\epsilon_{\phi}(x_{t}^{i},t,c,y))\frac{\partial g(\theta_{i},c)}{\partial\theta_{i}}\right],$ (2)
@@
$\min_{\phi}\mathbb{E}_{t,\epsilon,c}\left[\|\omega(t)(\epsilon_{\phi}(x_{t}^{i},t,c,y)-\epsilon)\|^{2}_{2}\right].$ (3)
@@
4 Our Approach
4.1 Analysis
Figure 2: We present a simulation of SDS (first row) and VSD (second row) in KL form on a 1D toy dataset, where the ground truth distribution $p_{\text{SD}}(x_{t}|y)$ is a 7-component Gaussian mixture model. Results are shown at $t=1$ (low noise data). In the third row and forth row, varying $p_{\text{SD}}(x_{t}|y^{\prime})$ with a new text prompt $y^{\prime}$ leads to diverse outcomes across different runs with SDS/VSD loss, motivating our approach.
Let us first motivate our method by an empirical analysis on the diversity of SDS and VSD loss, the prevailing loss functions for generating 3D assets from a given text prompt using a text-to-image model as prior. A notable drawback of this technique is that the SDS loss often yields almost identical results across different runs, primarily due to the mode-seeking behavior exhibited by the KL divergence between a Gaussian distribution and a multi-modal landscape of the text-to-image prior. More precisely, as shown from [38], the SDS loss in the KL form is given by:
$\mathcal{L}_{\text{SDS}}\triangleq\mathbb{E}_{t}\left[\text{KL}(q(x_{t}|x=g(\theta,c))||p_{\text{SD}}(x_{t}|y))\right].$ (4)
[…]
Meanwhile, the KL form of VSD (Eq. 5) is nearly identical to SDS, albeit the substitution of the unimodal Gaussian $q(x_{t}|x)$ with a more intricate, implicit distribution $q^{\mu}(x_{t}|c,y)=\mathbb{E}_{\mu(\theta|y)}[q(x_{t}|x=g(\theta,c))]$.
$\mathcal{L}_{\text{vsd}}\triangleq\mathbb{E}_{t}\left[\text{KL}(q^{\mu}(x_{t}|c,y)||p_{\text{SD}}(x_{t}|y))\right]$ (5)
By increasing the complexity of $q^{\mu}$, the optimized distribution $q^{\mu*}$ will possess a greater capacity to accurately fit the target distribution $p_{SD}(x_{t}|y)$. The VSD algorithm aims to draw data from $\mu^{*}(\theta|y)$ to minimize Eq. 5 through particle-based variational inference. As a result, the final 3D assets will exhibit a greater degree of variety.
We elucidate our intuition in Fig. 2 by performing a simulation for SDS and VSD on a 1D toy dataset, where $p_{\text{SD}}(x_{t}|y)$ is a Gaussian mixture model (GMM) comprising 7 components. We select a random position $x$ as the optimizable parameter for SDS, while the means and standard deviations of a 3-component GMM $\mu(\theta|y)$ serve as the optimizable parameters for VSD. Based on this analysis, instead of modeling $q(x_{t}|x)$ as done in SDS and VSD, we propose a new way to increase the diversity of the text-to-3D models by modifying the distribution $p_{\text{SD}}(x_{t}|y)$. Our idea is to diversify the condition $y$ to $y^{\prime}$ so that optimization with the prior $p_{\text{SD}}(x_{t}|y^{\prime})$ potentially yields a more substantial impact due to the alteration in the optimization landscape.
4.2 Method Overview
To implement the idea of diversifying $y$, inspired by textual inversion methods [11, 14] for 2D object personalization, we aim to condition text-to-3D synthesis such that the per-particle difference in $\epsilon_{\text{SD}}(x_{t}^{i},t,c,y^{\prime}_{i})$ is boosted by using different and distinct prompt $y^{\prime}_{i}$ for each particle $\theta_{i}$. To this end, we devise a new approach that leverages HiPer textual inversion [14] to enhance the resulting diversity. Here we base our discussion on VSD, but the idea generalizes to SDS as well.
Our approach consists of two stages: HiPer tokens inversion and textual score distillation. Firstly, we select a reference image $x^{r}_{i}$ for each particle and determine an optimized HiPer token $h^{*}_{i}$. This token is chosen such that the reference image can be faithfully reconstructed by the pretrained text-to-image diffusion model with the prompt $[y;h^{*}_{i}]$, where [;][;][ ; ] denotes concatenation. In the second stage, multiple particles $\theta_{i}$ are collectively optimized alongside a new shared domain adapter $\phi$, which is also encoded as a learnable token. This process forms the augmented text prompt $[y;h^{*}_{i};\phi]$ for each particle. The algorithm is depicted in Alg. 1 and illustrated in Fig. 3.
@@ 4.2 Method Overview
3:     Optimize $h_{i}$ given $x^{r}_{i}$ following  Eq. 6 to obtain $h_{i}^{*}$.
@@
5:     Update $\theta_{i}$ following Eq. 8.
6:     Update $\phi$ following Eq. 9.
@@ 4.3 HiPer tokens inversion
Figure 4: Optimization progress of VSD (upper) vs ours (lower). TSD with less #learnable parameters converges faster than VSD. Prompt: “A high-quality ice cream sundae”.
@@ 4.4 Textual score distillation (TSD)
$\nabla_{\theta_{i}}\mathcal{L}_{\text{TSD}}\triangleq\mathbb{E}_{t,\epsilon,c}\left[\omega(t)(\epsilon_{\text{SD}}(x_{t}^{i},t,y_{i})-\epsilon_{\text{SD}}(x_{t}^{i},t,y^{\prime}_{i}))\frac{\partial g(\theta_{i},c)}{\partial\theta_{i}}\right].$ (8)
@@
In Fig. 3 (Right), we show how we train the sampled NeRF model $\theta_{i}$ and shared token $\phi$ using the proposed TSD (Eq. 8) and MSE (Eq. 9) losses respectively. As can be seen in the Fig. 4, our method can produce higher quality samples than those produced by VSD [57] given the same amount of optimization time.
@@ 5 Experiments
Metrics. The common metrics for generative models such as FID [15] do not separately measure fidelity and diversity.
Inspired by [56], we proposed to use a modified version of Inception Quality (IQ) and Inception Variance (IV) to measure the quality and diversity of our models. Our IQ and IV are formulated as follows:
@@
We also propose the Cosine Sim metric to quantify the diversity of our particles. Specifically, for a given set of $K$ particles, we render the same view for each particle. The rendered images are then fed through a feature extractor (e.g., as DINO [63]) to obtain feature vectors. We then calculate the cosine similarity for these feature vectors across $\binom{K}{2}$ pairs and take the average. To ensure robustness, we also average the results over 120 rendered views.
Implementation details. Our method is implemented with threestudio, an open-source framework for text-to-3D synthesis [13]. We use $K=6$ particles for both VSD and our framework, TSD. The training of all experiments is conducted for 50K iterations. We use $L_{2}=5$ as HiPer tokens optimized in 1.4K iterations. Also, we use $L_{3}=8$ as shared learnable tokens $\phi$ and trained in 50K iterations. For the camera embedding, we use the same implementation as threestudio [13]. Regarding the resolution, we train each particle at $256\times 256$ for all methods.
@@
Table 1: Comparison with SOTAs
IQ $\downarrow$ IV $\uparrow$ Cosine Sim $\downarrow$
SDS [38] 3.695 4.577 0.720
VSD [57] 3.345 4.586 0.476
TSD (ours) 3.614 5.075 0.380
Table 2: Study on #HiPer tokens.
#tokens IQ $\downarrow$ IV $\uparrow$ Cosine Sim $\downarrow$
1 3.375 4.908 0.403
5 3.790 4.886 0.415
10 4.428 5.138 0.409
15 4.721 4.862 0.445
20 5.193 4.978 0.425
5.1 Comparison with Prior Methods
Baselines. We compare our method with two prominent text-to-3D methods including DreamFusion [38] and ProlificDreamer [57]. Note that we do not compare to other variants such as Magic3D [28] or Fantasia3D [8] since these methods address different issues of SDS, which is orthogonal to our method which focuses on diversity. For validation, we select a set of 60 text prompts from DreamFusion [38] and 10 text prompts generated randomly from ChatGPT [37].
Figure 6: We demonstrate that our method remains effective for SDS.
Figure 7: Visual results with different numbers of HiPer tokens.
Table 3: Ablation study of our method.
$\epsilon_{\phi}$ HiPer IQ $\downarrow$ IV $\uparrow$ Cosine Sim $\downarrow$ Training time (hours)
LoRA No 3.345 4.586 0.476 10.23
LoRA Yes 3.662 5.109 0.355 9.50
Shared tokens Yes 3.614 5.075 0.380 7.16